Wildhorse Company has net income of $1,230,700, assets of $12,188,200, and retains 69 percent of its income every year. What is the company’s internal growth rate? (Round answer to 1 decimal place, e.g. 15.5%.)

Internal Growth Rate %
Answer #1

ROA =Net Income/Total assets =1230700/12188200 =10.0975%
Retention ratio (b) =69%
Internal Growth Rate =ROA*b/(1-ROA*b) =10.0975% *69%/(1-10.0975% *69%) =7.49%or 7.50%